OpenStudy (clamin):

Which equation describes a parabola that opens left or right and whose vertex is at the point (h,v)? a. y=a(x-v)^2+h b. x=a(y-v)^2+h c. x=a(y-h)^2+v d. y=a(x-h)^2+v

OpenStudy (anonymous):

are you still there @clamin

OpenStudy (shouborno):

 Since the vertex is \( (h,v) \) , the equation is, \( x-h = a(y-v)^2\;, \;\;\;or,\;\;\;y-v=a(x-h)^2\) Now, if we put \(y=v+1,v-1\) in the equation \((b)\), we get the same value of \(x\), that is, \(x=a+h\). On the other hand, let's put \(x=h+1,h-1\) in the latter equation. Similar to the first equation, we get \(y=a+v \). \(\therefore \) We deduce, \((b)\) opens either right or left, according to the graph plotted above. \(QED\)